\begin{thebibliography}{99}

%\setlength{\parsep }{-0.5ex}
%\setlength{\itemsep}{-0.5ex}

%\frenchspacing

\newcommand\BAMS{\emph{Bull. Amer. Math. Soc.\ }}
\newcommand\BIT{\emph{BIT\ }}
\newcommand\Com{\emph{Computing\ }}
\newcommand\CA{\emph{Constr. Approx.\ }}
\newcommand\FCM{\emph{Found. Comput. Math.\ }}
\newcommand\JAT{\emph{J. Approx. Th.\ }}
\newcommand\JC{\emph{J. Complexity\ }}
\newcommand\JMA{\emph{SIAM J. Math. Anal.\ }}
\newcommand\JMAA{\emph{J. Math. Anal. Appl.\ }}
\newcommand\JMM{\emph{J. Math. Mech.\ }}
\newcommand\MC{\emph{Math. Comp.\ }}
\newcommand\NM{\emph{Numer. Math.\ }}
\newcommand\RMJ{\emph{Rocky Mt. J. Math.\ }}
\newcommand\SJNA{\emph{SIAM J. Numer. Anal.\ }}
\newcommand\SR{\emph{SIAM Rev.\ }}
\newcommand\TAMS{\emph{Trans. Amer. Math. Soc.\ }}
\newcommand\TOMS{\emph{ACM Trans. Math. Software\ }}
\newcommand\USSR{\emph{USSR Comput. Maths. Math. Phys.\ }}

\frenchspacing

\addcontentsline{toc}{chapter}{Bibliography}

%\bibitem{Ai11}
%C. Aistleitner, \emph{Covering numbers, dyadic chaining and discrepancy}, 
%J. Complexity 27, 531--540, 2011.
%
\bibitem{AHR15}
C. Aistleitner, A. Hinrichs, D. Rudolf,
On the size of the largest empty box amidst a point set, 
\emph{Discrete Applied Mathematics} {\bf 230} (2017), 146--150.
%%
\bibitem{BDDG14}
M. Bachmayr, W. Dahmen, R. DeVore and L. Grasedyck, 
Approximation of High-Dimensional Rank One Tensors, 
\emph{Constructive Approximation} {\bf 39}(2), 385--395, 2014.
%
%\bibitem{Ba59}       
%N. S. Bakhvalov,        
%On approximate computation of integrals,        
%\emph{Vestnik MGU, Ser. Math. Mech. Astron. Phys. Chem.}        
%{\bf 4}, 3--18, 1959, in Russian.       
       %
%% \bibitem{Ba71}             
%% N. S. Bakhvalov,        
%% On the optimality of linear methods for operator        
%% approximation in convex classes of functions.        
%% \USSR  {\bf 11}, 244--249, 1971.        
       %
%\bibitem{Ba14}
%Y. Baraud, \emph{Bounding the Expectation of the Supremum of an Empirical Process Over 
%a (Weak) VC-Major Class}, arXiv:1411.5571, 2014.

%\bibitem{BNR98}       
%V. Barthelmann, E. Novak and K. Ritter,       
%High dimensional polynomial interpolation on sparse grids,        
%\emph{Adv. in Comput. Math.} {\bf 12}, 273--288, 1999.        
%
%\bibitem{BP11}
%H. Brass and K. Petras,
%\emph{Quadrature Theory},
%American Math. Soc., Providence, 2011. 
%
%\bibitem{BG04}       
%H.-J.   Bungartz and  M. Griebel,        
%Sparse grids,        
%\emph{Acta Numerica} {\bf 13}, 147--269, 2004.        

\bibitem{Ca97}
J.W.S. Cassels, 
\emph{An introduction to the geometry of numbers}. 
Corrected reprint of the 1971 edition. Classics in Mathematics. 
Springer-Verlag, Berlin, 1997.

%%  \bibitem{CNR99}       
%%  R. Cools, E. Novak and K. Ritter,        
%%  Smolyak's construction of cubature formulas of arbitrary   
%%  trigonometric degree,        
%%  \Com   {\bf 62}, 147--162. 1999.        
%
%\bibitem{Dav67}
%P. J. Davis, 
%A construction of nonnegative approximate quadratures, 
%\MC {\bf 21}, 578--582, 1967.
%
%\bibitem{DKS}
%J. Dick, F. Y. Kuo and I. H. Sloan, 
%High-dimensional integration: The quasi-Monte Carlo way,
%\emph{Acta Numerica} {\bf 22}, 133--288, 2013.
%
%\bibitem{DP10} 
%J. Dick and F. Pillichshammer,
%\emph{Digital nets and sequences}, 
%Cambridge University Press, Cambridge, 2010.
%%
%\bibitem{DP14} 
%J. Dick and F. Pillichshammer,
%\emph{Discrepancy theory and quasi-Monte Carlo integration}, 
%A panorama in Discrepancy Theory, Lecture Notes in Math. 2107, 
%Springer Verlag, 2014.
%%
%\bibitem{DT97}
%M. Drmota, R. F. Tichy, Sequences, discrepancies and applications.
%Springer Lecture Notes in Math. 1651, 1997.
%
%\bibitem{Du} V.V. Dubinin, Cubature formulas for classes of functions
%with bounded mixed difference, Matem. Sbornik, 183 (1992), issue 7;
 %English transl. in  Math.
%USSR Sbornik 76 (1993),  283--292.
%
%\bibitem{Du2} V.V. Dubinin, Cubature formulae for Besov classes,
%Izvestiya Math 61(2) (1997), 259--83.
%
%\bibitem{Dudley99}
%R. M. Dudley,
%\emph{Uniform central limit theorems},
%Cambridge University Press, Cambridge, 1999.
%
\bibitem{DJ13}
A. Dumitrescu and M. Jiang, 
On the largest empty axis-parallel box amidst n points, 
\emph{Algorithmica} 66(2):225--248, 2013.

\bibitem{DJ13b}
A. Dumitrescu and M. Jiang, 
Maximal empty boxes amidst random points, 
\emph{Combin. Probab. Comput.} {\bf 22}(4), 477--498, 2013.

%\bibitem{DU14}
%D. D\~ung and T. Ullrich, 
%Lower bounds for the integration error for multivariate functions with mixed smoothness 
%and optimal Fibonacci cubature for functions on the square, 
%\emph{Math. Nachrichten}, DOI 10.1002/mana.201400048, 2014.

%
%\bibitem{FR97}
%B. Fine and G. Rosenberger,
%\emph{The fundamental theorem of algebra}, 
%Undergraduate Texts in Mathematics, Springer-Verlag, New York, 1997.

%\bibitem{FeSt71} Fefferman, C. and Stein, E. M., Some maximal inequalities, \emph{Amer. J. Math.} $\bf 93$,
%107--115, 1971. 

%\bibitem{FrJa90}
%M.~{F}razier and B.~{J}awerth.
%\newblock {A} discrete transform and decompositions of distribution spaces.
%\newblock {\em Journ. of Funct. Anal.}, 93:34--170, 1990.
%
\bibitem{Fr76}
K. K. Frolov,
Upper error bounds for quadrature formulas on function classes, 
\emph{Dokl. Akad. Nauk SSSR} 231, 818--821, 1976.
%

%\bibitem{Fr77}
%K. K. Frolov,
%The connection of quadrature formulas and sublattices of the 
              %lattice of integer vectors, 
%\emph{Dokl. Akad. Nauk SSSR} {\bf 232}, 40--43, 1977.
%
\bibitem{Fr80}
K. K. Frolov,
Upper bound of the discrepancy in metric {$L_{p}$}, {$2\leq p<\infty $}, 
\emph{Dokl. Akad. Nauk SSSR} {\bf 252}, 805--807, 1980.
    %
%\bibitem{GG98}       
%Th. Gerstner and  M. Griebel,        
%Numerical integration using sparse grids,        
%\emph{Num. Algorithms} {\bf 18}, 209--232, 1998.        
       %
%\bibitem{glasserman2004monte}
%P.~Glasserman.
%\newblock {\em Monte Carlo Methods in Financial Engineering}.
%\newblock Applications of mathematics : stochastic modelling and applied
  %probability. Springer, 2004.
%
%
%\bibitem{Gr06}       
% M. Griebel,        
%Sparse grids and related approximation schemes for higher        
%dimensional problems,        
%in:  \emph{Foundations of Computational Mathematics},        
%Santander 2005,        
%L. M. Pardo et al. (eds.), 106--161,       
%London Math. Soc., Lecture Note Series 331,        
%Cambridge University Press, 2006.   
%
\bibitem{GL87}
P.M. Gruber, C.G. Lekkerkerker, 
\emph{Geometry of Numbers}.
Second edition. North-Holland Mathematical Library, 37. 
North-Holland Publishing Co., Amsterdam, 1987.
%
%\bibitem{HNWW01}
%S. Heinrich, E. Novak, G. W. Wasilkowski, H. Wo\'zniakowski,
%The inverse of the star-discrepancy depends linearly on the dimension. 
%\emph{Acta Arithmetica} {\bf 96}, 279--302, 2001.
%
%\bibitem{HNUW12}
%A. Hinrichs, E. Novak, M. Ullrich, H. Wo\'zniakowski,
%The curse of dimensionality for numerical integration
%of smooth functions,
%\emph{Math. Comp.}, to appear. 
%

%\bibitem{Hi04}
%A.~Hinrichs,
%\emph{Covering numbers, Vapnik-\v{C}ervonenkis classes and bounds for the star-discrepancy}, 
%J. Complexity 20(4):477--483, 2004.

%\bibitem{Hi10}
%A.~Hinrichs.
%\newblock Discrepancy of {H}ammersley points in {B}esov spaces of dominating
  %mixed smoothness.
%\newblock {\em Math. Nachr.}, 283(3):478--488, 2010.
%
%\bibitem{HiMaOeUl14}
%A. Hinrichs, L. Markhasin, J. Oettershagen, T. Ullrich, 
%Optimal quasi-Monte Carlo rules on higher order digital nets for the numerical integration of multivariate periodic functions,
%preprint Bonn 2014. 
%
%\bibitem{Hl62}
%E.~Hlawka,
%\newblock Zur angen\"aherten {B}erechnung mehrfacher {I}ntegrale.
%\newblock {\em Monatsh. Math.}, 66:140--151, 1962.
%
%\bibitem{Hl76}
%E.~Hlawka,
%\emph{Absch\"atzung von trigonometrischen {S}ummen mittels diophantischer {A}pproximationen}, 
%\"Osterreich. Akad. Wiss. Math.-Naturwiss. Kl. S.-B. II, 185:43--50, 1976.
%
%\bibitem{HNUW14}
%A. Hinrichs, E. Novak, M. Ullrich, H. Wo\'zniakowski,
%The curse of dimensionality for numerical integration
%of smooth functions II,
%\emph{J. Complexity} 30, 117-143, 2014.
%
%\bibitem{HO14}
%A. Hinrichs, J. Oettershagen, 
%Optimal point sets for quasi–-Monte Carlo integration of bivariate periodic 
%functions with bounded mixed derivatives, arXiv:1409.5894, 2014.

%\bibitem{HZ07}       
%F. L. Huang and  S. Zhang,        
%Approximation of infinitely differentiable multivariate        
%functions is not strongly tractable,       
%\JC  {\bf 23}, 73--81, 2007.        
       %
\bibitem{Kr17}
D. Krieg,
On the Dispersion of Sparse Grids,
preprint, arXiv:1709.02983, 2017.

\bibitem{KN16}
D. Krieg and E. Novak, A Universal Algorithm for Multivariate Integration, 
\emph{Found. Comput. Math.} {\bf 17} (2016), no. 4, 895--916.
%DOI:10.1007/s10208-016-9307-y.

%% \bibitem{KW05}     
%% A. Klimke and B. Wohlmuth,      
%% Algorithm 847: spinterp: piecewise multilinear hierarchical      
%% sparse grid interpolation in MATLAB ,     
%% \emph{ACM Trans. Math. Software} {\bf 31}, 561--579, 2005.      
%
%\bibitem{LaPC}
%G. Larcher, \emph{personal communication}, 2015.

%\bibitem{LeeW11}
%C.-L. Lee and K. B. Wong,
%On Chebyshev's polynomials and certain combinatorial identities,
%\emph{Bull. Malays. Math. Sci. Soc. (2)} {\bf 34}, 279--286, 2011.

%\bibitem{Ly03}
%J. N. Lyness,
%Notes on Lattice Rules,
%\JC {\bf 19}, 321--331, 2003.
%
%
 %\bibitem{Ma13}
%L. Markhasin, Discrepancy and integration in function spaces with dominating 
%mixed smoothness, \emph{Dissertationes Math.} 494, 1--81, 2013.

%
%\bibitem{MG}
%Th. M\"uller-Gronbach,
%Hyperbolic cross designs for approximation 
%of random fields,
%\emph{J. Statistical Planning and Inference} {\bf 66}, 
%321--344, 1998.
%
\bibitem{NLH84}
A. Naamad, D. T. Lee, and W.-L. Hsu, 
On the maximum empty rectangle problem, 
\emph{Discrete Appl. Math.} 8(3), 267–277, 1984.
%
\bibitem{NUU16}
V. K. Nguyen, M. Ullrich and T. Ullrich, 
Change of variable in spaces of mixed smoothness and numerical integration of 
multivariate functions on the unit cube, 
\emph{Constr. Approx.} {\bf 46} (2017), 69--108.
%
\bibitem{Niederreiter83}       
H. Niederreiter,        
A quasi-{M}onte {C}arlo method for the approximate computation of the 
extreme values of a function,       
\emph{Studies in Pure Mathematics}, pp.~523--529, Birkh\"auser, Basel, 1983.
%
%\bibitem{Niederreiter92}       
%H. Niederreiter,        
%\emph{Random Number Generation and quasi-Monte Carlo Methods},       
%Society for Industrial and Applied Mathematics, Philadelphia, 1992.
%
%\bibitem{No88}       
%E. Novak,        
%\emph{Deterministic and Stochastic Error Bounds in      
%Numerical Analysis},       
%LNiM {\bf 1349}, Springer Verlag, Berlin, 1988.   
%     
\bibitem{No15}       
E. Novak,        
\emph{Some Results on the Complexity of Numerical Integration},       
arXiv:1409.6714, 2015.       
%
%\bibitem{NR96}       
%E. Novak and K. Ritter,        
%High dimensional integration of       
%smooth functions over cubes,        
%\NM   {\bf 75}, 79--97, 1996.        
       %
%\bibitem{NR97}       
%E. Novak and  K. Ritter, The curse of dimension and a       
%universal method  for numerical integration,        
%in:   \emph{Multivariate Approximation and Splines},         
%G. N\"urnberger, J. W. Schmidt and       
%G. Walz, eds.,  ISNM {\bf 125}, Birkh\"auser, Basel,  177--188, 1997.        
       %
%\bibitem{NR99}       
%E. Novak and K. Ritter,        
%Simple cubature formulas with high polynomial exactness,        
%\CA   {\bf 15}, 499--522, 1999.        
       %
%\bibitem{NRSS99}       
%E.   Novak, K.  Ritter, R.  Schmitt and  A. Steinbauer,        
%On an interpolatory  method for high dimensional integration,       
%\emph{J. Comput. Appl. Math.} {\bf 112}, 215--228, 1999.        
%
\bibitem{NR15}
E. Novak, D. Rudolf, 
Tractability of the approximation of high-dimensional rank one tensors, 
\emph{Constructive Approximation}, doi:10.1007/s00365-015-9282-6, 2015.
%
%\bibitem{NW08}       
%E. Novak and H. Wo\'zniakowski,        
%\emph{Tractability of Multivariate Problems},        
%Volume 1, Linear Information, European Math. Soc., Z\"urich,        
%2008.        
%
%\bibitem{NW09b}    
%E. Novak and H. Wo\'zniakowski,    
%Approximation of infinitely differentiable multivariate functions    
%is intractable,    
%\JC {\bf 25}, 398--404, 2009. 
%
\bibitem{NW10}      
E. Novak and H. Wo\'zniakowski,             
\emph{Tractability of Multivariate Problems},      
Volume II: Standard Information for Functionals,      
European Math. Soc. Publ. House, Z\"urich,              
2010.       
%
%%  \bibitem{NW12}      
%%  E. Novak and H. Wo\'zniakowski,             
%%  \emph{Tractability of Multivariate Problems},      
%%  Volume III: Standard Information for Operators, 
%%  European Math. Soc. Publ. House, Z\"urich,              
%%  2012.       
%
%\bibitem{Pe94}
%A. P. Petukhov,
%Convergence of Fourier series for functions in the classes of Besov-Lizorkin-Triebel,
%\emph{Mathematical Notes} {\bf 56}, 694--698, 1994.

%\bibitem{RT96}
%G. Rote and R. F. Tichy, 
%\emph{Quasi-Monte Carlo methods and the dispersion of point sequences}, 
%Math. Comput. Modelling 23(8-9):9--23, 1996.
%
%\bibitem{Ro54}
%K. F. Roth, \emph{On irregularities of distribution}, Mathematika 1:73--79, 1954.
%
%\bibitem{Ru17}
%D. Rudolf, 
%An Upper Bound of the Minimal Dispersion via Delta Covers, 
%preprint, arXiv:1701.06430, 2017.
 
 
\bibitem{Sk94}
M.M. Skriganov, 
Constructions of uniform distributions in terms of geometry of numbers, 
\emph{Algebra i Analiz} 6, 200--230, 1994.


%\bibitem{SW71}
%E. M. Stein and G. Weiss,
%\emph{Introduction to {F}ourier analysis on {E}uclidean spaces}, 
%Princeton Mathematical Series, No. 32, 
%Princeton University Press, Princeton, N.J., 1971.


%\bibitem{Sm63}        
%S. A. Smolyak,       
%Quadrature and interpolation formulas for tensor products       
%of certain classes of functions,       
%\emph{Dokl. Akad. Nauk SSSR} {\bf 4}, 240--243, 1963.        
%
\bibitem{Sos17}
J. Sosnovec, 
A note on the minimal dispersion of point sets in the unit cube, 
preprint, arXiv:1707.08794, 2017.

\bibitem{TW16}
N. Technau, M. Widmer, 
On a counting theorem of Skriganov, 
preprint, arXiv:1611.02649v1, 2016.
%
\bibitem{Te93}
V. N. Temlyakov, 
\emph{Approximation of periodic functions}, 
Computational Mathematics and Analysis Series, 
Nova Science Publishers, Inc., Commack, NY, 1993.
%
\bibitem{Te03}
V. N. Temlyakov, 
Cubature formulas, discrepancy, and nonlinear approximation, 
\JC 19, 352--391, 2003.

\bibitem{Te17a}
V. N. Temlyakov, 
The Marcinkewiecz-type discretization theorems for
the hyperbolic cross polynomials, 
arXiv:1702.01617v2, 2017.

\bibitem{Te17b}
V.N. Temlyakov, 
The Marcinkiewicz-type discretization theorems,
preprint, arXiv:1703.03743v1, 2017.

\bibitem{Te17c}
V. N. Temlyakov,
Universal discretization,
preprint, arXiv:1708.08544, 2017.

\bibitem{Te17d}
V.N. Temlyakov, 
Dispersion of the Fibonacci and the Frolov point sets,
preprint, arXiv:1709.08158, 2017.
%
\bibitem{MU18}
M. Ullrich, 
A lower bound for the dispersion on the torus, 
\emph{Math. Comput. Simulation} {\bf 143}, 186--190, 2018.

\bibitem{MU15}
M. Ullrich,
On ``Upper error bounds for quadrature formulas on function classes'' by K. K. Frolov. 
In: Ronald Cools and Dirk Nuyens (Eds): 
\emph{Monte Carlo and Quasi-Monte Carlo Methods}, 
Springer Proceedings in Mathematics \& Statistics {\bf 163}, 
571--582, 2016.

\bibitem{MU17}
M. Ullrich, 
A Monte Carlo method for integration of multivariate smooth functions, 
\emph{SIAM J. Numer. Anal.} {\bf 55}(3), 1188-1200, 2017.

\bibitem{UU16}
M. Ullrich and T. Ullrich, The Role of Frolov's Cubature Formula for Functions with Bounded Mixed Derivative, 
\emph{SIAM J. Numer. Anal.} {\bf 54}(2), 969--993, 2016.

\bibitem{UV17}
M. Ullrich and J. Vyb\'iral, 
An upper bound on the minimal dispersion, 
preprint, 2017.

%\bibitem{Vyb06}
%J. Vyb\'iral,
%Function spaces with dominating mixed smoothness,
%\emph{Dissertationes Math.} 436, 73 pp., 2006.


%\bibitem{YLV00}
%S. Yakowitz, P. L'Ecuyer and F. V{\'a}zquez-Abad,
%\emph{Global stochastic optimization with low-dispersion point sets}, 
%Oper. Res. 48(6):939--950, 2000.
							
\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{latexsym,amsfonts,amsthm,amssymb, epsfig}
\newtheorem{thm}{Theorem}

\newtheorem{prop}[thm]{Proposition}
\newtheorem{cor}[thm]{Corollary}

\theoremstyle{remark}
\newtheorem{rem}{Remark}

\theoremstyle{definition}

\newcommand{\R}{\mathbb{R}}

\newcommand{\N}{\mathbb{N}}

\newcommand{\Z}{\mathbb{Z}}

\newcommand{\calP}{\mathcal{P}}

\newcommand{\LL}{\mathbb{L}}

\DeclareMathOperator{\disp}{disp}

\title[Dispersion of admissible lattices]{A note on the dispersion\\ of admissible lattices}
\author{Mario Ullrich}
\date{\today}

\begin{document}

\begin{abstract}
In this note we show that the volume of axis-parallel boxes in $\R^d$ which do not 
intersect an admissible lattice $\LL\subset\R^d$ is uniformly bounded. 
In particular, this implies that the dispersion of the dilated lattices $N^{-1/d}\LL$ 
restricted to the unit cube is of the (optimal) order $N^{-1}$ as $N$ goes to infinity.
This result was obtained independently by V.N.~Temlyakov (arXiv:1709.08158).
\end{abstract}

\maketitle

Let $\LL\subset\R^d$, $d\in\N$, be an \emph{admissible lattice}
(in the sense of Skriganov~\cite{Sk94}), 
i.e.,  $\LL=T(\Z^d)$ for some $T\in\R^{d\times d}$ such that
\begin{equation}\label{eq:adm}
{\rm Nm}(\LL) \,:=\, \inf_{z\in\LL\setminus\{0\}}\, \prod_{j=1}^d |z_j| \,>\, 0.
\end{equation}
Such lattices play a crucial role in the geometry of numbers, see e.g.~\cite{Ca97,GL87} 
and the references therein, 
especially in the theory of Diophantine approximation.
Moreover, they are among the most important explicit constructions of 
point sets, which satisfy various \emph{uniform distribution properties}, like the 
optimal order of the discrepancy~\cite{Fr80,Sk94}.
%Only recently, such lattices 
They attracted quite a lot of attention in numerical analysis 
as the corresponding point sets seem to serve as an optimal and universal choice as 
nodes for corresponding cubature rules, 
see e.g.~\cite{Fr76,KN16,NUU16,Sk94,Te93,Te03,MU15,MU17,UU16}.
See also \cite{No15,NW10,Te03} for surveys on the state of the art in numerical integration.

Apart from the concepts and geometric quantities that are important for the above, 
there is an increasing interest in the dispersion of a point set. 
For $d\in\N$ and a point set $\calP\subset[0,1]^d$, 
the \emph{dispersion of $\calP$} is defined by
\[
\disp(\calP) \;:=\; \sup_{B\colon B\cap\calP=\varnothing}\, |B|, 
\]
where the supremum is over all axis-parallel boxes $B=I_1\times\dots\times I_d$ 
with intervals $I_\ell\subset[0,1]$, 
and $|B|$ denotes the volume of $B$.
This quantity was proven to be essential for various numerical problems 
including optimization~\cite{Niederreiter83}, approximation of high-dimensional 
rank-1 tensors~\cite{BDDG14,NR15} and, very recently, in the study 
of Marcinkiewicz-type discretization theorems~\cite{Te17a,Te17b,Te17c}. 
Moreover, algorithms for finding such a box with maximal volume, 
with a motivation coming from information theory, are also of some interest, 
see~\cite{DJ13,DJ13b,NLH84} and references therein.
In all these contexts, it is desirable to have tight bounds on the dispersion 
(clearly depending on $d$ and $\#\calP$), and to find explicit constructions 
of point sets satisfying these bounds.
The best bounds so far, with special emphasis on the dependence on the dimension $d$, 
can be found in the recent articles~\cite{AHR15,Kr17,Sos17,MU18,UV17}. 
In particular, the only presently known optimal 
(w.r.t.~the dependence on $\#\calP$) 
construction is based on $(t,m,d)$-nets, see~\cite[Section~4]{AHR15}.
Here, we will prove that admissible lattices also lead to optimal 
(w.r.t.~the dependence on $\#\calP$) constructions for the dispersion.
This was obtained independently by Temlyakov~\cite{Te17d}. 
However, we show that this is a quite direct corollary of the famous 
lattice point counting result of Skriganov~\cite{Sk94}, which seems 
to be of independent interest. 
%In any case, our proof is much shorter.

In this note the point sets under consideration are admissible lattices $\LL$ 
and we consider the (apparently larger) \emph{lattice-dispersion of $\LL$}, 
i.e., the quantity
\[
\disp^*(\LL) \;:=\; \sup_{B\colon B\cap\LL=\varnothing}\, |B|,
\]
where the supremum is over all axis-parallel boxes $B=I_1\times\dots\times I_d$ 
with intervals $I_\ell\subset\R$, 
and $|B|$ denotes again the volume of $B$.

First, we prove the following result.

\begin{prop}\label{prop:main}
Let $\LL\subset\R^d$ be a lattice that satisfies~\eqref{eq:adm}.
Then, 
\[
\disp^*(\LL) \,\le\, D_\LL \,<\, \infty,
\]
where $D_\LL$ depends on $\LL$ only by means of 
$\det(\LL)$ and ${\rm Nm}(\LL)$.
\end{prop}
\medskip

\begin{rem}
In the application of such results, particularly in numerical analysis, 
it is often necessary to switch between properties of the lattice 
$\LL=T(\Z^d)$ and the corresponding \emph{dual lattice} $\LL^*=(T^{-1})^\top(\Z^d)$.
Let us therefore note that $\LL$ 
%is admissible (see~\eqref{eq:adm}) 
satisfies~\eqref{eq:adm}
if and only if 
$\LL^*$ satisfies~\eqref{eq:adm}, see~\cite[Lemma~3.1]{Sk94}. 
\end{rem}

\bigskip

From this result we obtain a bound on the dispersion (in the unit cube) 
for dilated admissible lattices. 
For this, we introduce the notation 
\[
t^{-1} \LL:=\{(t_1^{-1}z_1,\dots,t_d^{-1}z_d)\colon z\in\LL\}
\] 
and
$[0,t]:=[0,t_{1}]\times\dots\times[0,t_{d}]$
for $t=(t_1,\dots,t_d)\in\R^d$ with 
%and define
$n(t):=\bigl|[0,t]\bigr|=\prod_{j=1}^d|t_j|>0$.

\begin{cor}\label{cor1}
Let $\LL\subset\R^d$ be a lattice that satisfies~\eqref{eq:adm}.
Then, 
\[
\disp\Bigl((t^{-1} \LL)\cap[0,1]^d\Bigr) \,\le\, \frac{D_\LL}{n(t)}.
\]
Moreover, for every $N\in\N$ there exist $t_N\in\R^d$ such that
$\calP_N:=(t_N^{-1} \LL)\cap[0,1]^d$ satisfies $\#\calP_N=N$ and
\[
\disp(\calP_N) \,\le\, \frac{C_\LL}{N}
\]
for some constant $C_\LL$ that
%and $\widetilde{C}(\LL)$ 
depends on $\LL$ only by means of 
$\det(\LL)$ and ${\rm Nm}(\LL)$.
\end{cor}
\medskip

\begin{proof}[Proof of Proposition~\ref{prop:main}]
We know from the work of Skriganov~\cite[Theorem~1.1]{Sk94} that, 
for $B_{x,t}:=x+[0,t]$, $x,t\in\R^d$, we have
\[\begin{split}
\Bigl|\#\bigl(\LL\cap B_{x,t}\bigr) - \frac{|B_{x,t}|}{\det(\LL)}\Bigr|
\,\le\, D'_\LL\, \ln\Bigl(2+n(t)\Bigr)^{d-1}
\end{split}\]
uniformly in $x$, where the constant $D'_\LL$ 
depends on $\LL$ only by means of $\det(\LL)$ and ${\rm Nm}(\LL)$.
If we now assume that the (shifted and dilated) box $B_{x,t}$ 
contains no point of $\LL$, we obtain that
\[
|B_{x,t}| \,\le\, \det(\LL)\,D'_\LL\, \ln\Bigl(2+n(t)\Bigr)^{d-1}.
\]
This easily implies that $|B_{x,t}| = n(t) \le D_\LL$ for 
some $D_\LL<\infty$ that only depends on $\det(\LL)$ and $D'_\LL$.

%Assume to the contrary, $\disp^*(\LL)=\infty$ would imply
%$\disp^*(t^{-1}\LL)=\infty$ for any $t=(t_1,\dots,t_d)\in\R^d$.
%Here, we use the notation 
%$t^{-1} \LL=\{(t_1^{-1}z_1,\dots,t_d^{-1}z_d)\colon z\in\LL\}$.
%In particular, this would imply the existence of 
%$x\in\R^d$ and a sequence of $t=(t_1,\dots,t_d)\in\R^d$ with 
%$n(t):=\prod_{j=1}^d|t_j|\to\infty$, 
%such that $\#\bigl((x+t^{-1} \LL) \cap[0,1]^d\bigr)=0$ and 
%therefore
%\[
%\Bigl|\frac{\#\bigl((x+t^{-1} \LL) \cap[0,1]^d\bigr)}{n(t)}-\vol_d\bigl([0,1]^d\bigr)\Bigr| 
%\,=\, 1.
%\]
%However, we know from the work of Skriganov~\cite{Sk94} that 
%\[
%\Bigl|\frac{\#\bigl((x+t^{-1} \LL) \cap[0,1]^d\bigr)}{n(t)}-\vol_d\bigl([0,1]^d\bigr)\Bigr| 
%\,\le\, C\, \frac{\log^{d-1}(n(t))}{n(t)},
%\]
%uniformly in $x$ and the chosen sequence.
%The right hand side goes to zero as $n(t)\to\infty$: a contradiction.
\end{proof}

\medskip

\begin{proof}[Proof of Corollary~\ref{cor1}]
We obtain from Proposition~\ref{prop:main} that under the given assumption 
$\disp^*(\LL)\le D_\LL<\infty$. Therefore, we obtain from the definitions 
and the homogeneity of the lattice-dispersion that
\[
\disp\Bigl((t^{-1} \LL)\cap[0,1]^d\Bigr)
\,\le\, \disp^*\Bigl(t^{-1} \LL\Bigr)
\,=\, \frac{\disp^*\bigl(\LL\bigr)}{n(t)}
\,\le\, \frac{D_\LL}{n(t)}.
\]

For the second part of the corollary, note that the existence of 
%an 
infinitly many 
$t_N$ with $\#\calP_N=N$ follows from the fact that every hyperplane 
$\{x\colon x_j=a\}$, $a\in\R$, 
contains at most one point of $\LL$, see e.g.~\cite[Remark~2.1]{Sk94}. 
Now note that also $N=\#\bigl(\LL\cap[0,t_N]\bigr)$, 
%where $[0,t_N]:=[0,t_{N,1}]\times\dots\times[0,t_{N,d}]$, 
and that \eqref{eq:adm} implies $n(t_N)>{\rm Nm}(\LL)$ for $N\ge2$.
We partition $[0,t_N]$ into at most 
\[
|[0,t_N]|/{\rm Nm}(\LL)+1=n(t_N)/{\rm Nm}(\LL)+1\le 2\, n(t_N)
\]
axis-parallel boxes with volume less than ${\rm Nm}(\LL)$. 
Each of these boxes contains at most one point from $\LL$.
Otherwise, two distinct points $x,y\in\LL$ in the box must satisfy 
$\prod_{j=1}^d|x_j-y_j|<{\rm Nm}(\LL)$: a contradiction to~\eqref{eq:adm}.
Therefore, we obtain $N\le 2\, n(t_N)$, which proves the result 
with $C_\LL=2 D_\LL$.\\
\end{proof}

\goodbreak

\end{document}